please show the equation for adding mixed numbers example 4/7+3/14
Answer by edjones(8007)

We have to find the least common denominator (lcd).
in this case we know that 7 goes into 14 twice so we can use 14 as the lcd.
4/7+3/14
4/7=8/14
So, 8/14+3/14=11/14 Answer
Ed